OpenStudy (anonymous):

Sumone help me setp by steeepppp ::)) ---Solve the system, using substitution. Write the solution as an ordered pair leaving no spaces, or write no solution or infinitely many solutions; Example: (2,1) 2x + 3y = 4 and x - 2y = -5

OpenStudy (amistre64):

what is you substitution value?

OpenStudy (anonymous):

3y?

OpenStudy (amistre64):

not quite; that x-2y=-5 looks like we can solve it for "x" the easiest

OpenStudy (amistre64):

what would a "value" for x be using that equation?

OpenStudy (anonymous):

2.58,1.25

OpenStudy (anonymous):

x=-7

OpenStudy (anonymous):

if someone is trying to explain and teach something, its a little rude to just throw out answers like that.

OpenStudy (amistre64):

the word "value" here is not to imply a solid value such as -7 or 13 or any number; but a value in terms of y x -2y = -5 ; lets add 2y to each side +2y +2y ---------- x = 2y-5 ; whatever y equals this is the value for x

OpenStudy (anonymous):

@Joemath Tru that! But everybdy is trying to help tho.

OpenStudy (anonymous):

3x+3y=4 becomes x=(4/3)-y substitude into equation 2 (4/3)-y-2y=-5 basic math y=1.25 substitude the answer into equation 1 x=2.58

OpenStudy (amistre64):

the process is more important the the numbers; the numbers always change while the process remains the same

OpenStudy (amistre64):

do you see how I got a value for x, in terms of y?

OpenStudy (anonymous):

@amister64 Yes .

OpenStudy (amistre64):

good, because that is the important step in substitutions; now we can use that "value" in the other equation and just have to solve for "y" alone

OpenStudy (amistre64):

x - 2y = -5 x = 2y-5 ------------------- 2x + 3y = 4 2(2y-5) + 3y = 4 ; the rest is just algebring it into submission :)

OpenStudy (anonymous):

Then you add like terms right???

OpenStudy (amistre64):

2(2y-5) + 3y = 4 ; distribute the 2 thru 4y -10 +3y = 4 ; combine like terms 7y -10 = 4 ; add 10 across +10 +10 ------------ 7y = 14 ; and divide off the 7 /7 /7 -------- y = 2

OpenStudy (amistre64):

now that we have a solid value for "y"; y = 2 , we can find a good solid value for x. recall that we determined a value for x as: x = 2y-5 , now we know y=2 so plug it in x = 2(2) - 5 x = 4-5 = -1

OpenStudy (amistre64):

its always good to include as a last step to dbl chk your work

OpenStudy (amistre64):

an easy way to check is simply: http://www.wolframalpha.com/input/?i=2x+%2B+3y+%3D+4+and+x+-+2y+%3D+-5 :)

OpenStudy (anonymous):

Thanks Alot : ))))

OpenStudy (amistre64):

youre welcome; hope it helped ;) so to wrap it up; your answer should be: (-1,2) i believe

OpenStudy (anonymous):

Ok